# The Spatial Hydrodynamic Attractor: Resurgence of the Gradient Expansion

### 23 Mar 2026 v1

@@ Abstract
Far-from-equilibrium kinetic systems collapse onto a hydrodynamic attractor, traditionally approximated by a gradient expansion. While temporal gradient series are non-Borel summable and require transseries completions, the analytic structure of the spatial expansion has remained elusive. Here, we derive exact closed-form Chapman–Enskog coefficients at all orders via Lagrange inversion and prove that the non-relativistic spatial gradient series, though factorially divergent, is strictly Borel summable. Furthermore, we show that this divergence originates from unbounded Galilean velocities; enforcing relativistic causality yields a convergent spatial hydrodynamic expansion with finite radius.
@@
The classical treatment of non-equilibrium systems in kinetic theory relies on a systematic gradient expansion of the distribution function, truncated at finite order to yield macroscopic fluid dynamics. The validity of this expansion is controlled by the Knudsen number, the ratio of microscopic kinetic scales to the macroscopic scales of the coarse-grained hydrodynamic description. In relativistic systems, gradients of macroscopic fields normalized by the local temperature play an analogous role. However, it is well established that the resulting gradient expansion is purely asymptotic and diverges factorially [5].
@@
Despite this formal divergence, numerical [6] and experimental [1] evidence has shown that far-from-equilibrium systems can still be quantitatively described by hydrodynamics, suggesting the existence of a hydrodynamic attractor. Therefore, the divergent nature of the gradient expansion does not rule out the existence of such a stable manifold.
@@
This phenomenon was first explored for temporal gradient expansions by Heller and Spaliński [10]. They demonstrated that for the highly symmetric, longitudinally expanding quark-gluon plasma (Bjorken flow) within Müller-Israel-Stewart framework, the fluid dynamics reduce to a set of ordinary differential equations. They showed that despite the factorial divergence of the hydrodynamic series at large orders, the attractor rigorously exists. Its exact reconstruction, requires generalized Borel resummation; the series is non-Borel summable along the positive real axis due to singularities corresponding to the decay of non-hydrodynamic modes, necessitating a full transseries completion.
@@
This decaying of non-hydrodynamic modes (attractor paradigm) was subsequently generalized by Romatschke [14], who demonstrated that conformal Bjorken flow across three distinct microscopic theories: a variant of second-order BRSSS hydrodynamics [3], Boltzmann theory, and strongly coupled $\mathcal{N}=4$ supersymmetric Yang-Mills (via AdS/CFT), universally collapses onto a hydrodynamic attractor.
@@
These foundational studies focused exclusively on the temporal (longitudinal) gradient expansion, establishing that its divergence universally requires generalized Borel resummation. Unlike temporal evolution, spatial gradients trigger short-wavelength ultraviolet pathologies, such as the Burnett instability; the Borel summability and resurgent structure of the spatial gradient series has not been systematically studied.
@@
In this letter, we show that for non-relativistic Boltzmann kinetic theory, the spatial gradient expansion is factorially divergent but strictly Borel summable, unambiguously reconstructing the spatial attractor; moreover, once relativistic causality is imposed on the velocity space, the divergence is cured entirely, yielding a strictly convergent series with a finite radius.
@@
Starting from the one-dimensional kinetic Bhatnagar-Gross-Krook (BGK) equation [4] and setting the particle mass $m$, thermal energy $k_{B}T$, and relaxation time $\tau$ to unity, we write the non-dimensionalized kinetic equation for the distribution function $f(x,v,t)$ as
@@
In Fourier space this means $\hat{Z}_{\text{re}}=\hat{\Theta}_{\text{re}}(\lambda,k^{2})\,\hat{n}$ and $\hat{Z}_{\text{im}}=ik\,\hat{\Theta}_{\text{im}}(\lambda,k^{2})\,\hat{n}$, where the factor $ik$ in $\hat{Z}_{\text{im}}$ reflects the odd parity of the flux. The central requirement is the dynamic invariance condition: the macroscopic and microscopic time derivatives must agree exactly on this manifold [9, 11], bridging the kinetic and hydrodynamic descriptions in the spirit of Hilbert’s sixth problem.
@@
For small $k$ (equivalently large $A$), the resolvent is expanded as a geometric series, $\frac{A}{A+v^{2}}=\sum_{m=0}^{\infty}(-1)^{m}(v^{2}/A)^{m}$, and integrated term by term using the Gaussian moments $\langle v^{2m}\rangle=(2m{-}1)!!$ :
@@
where $x=k^{2}(1{+}\hat{\omega})^{-2}=A^{-1}$. Since $x$ itself depends on $\hat{\omega}$, Eq. (Exact Chapman-Enskog Coefficients.—) is an implicit relation for $\hat{\omega}(k)$. The Chapman–Enskog (CE) coefficients $\hat{\omega}=\sum_{n=1}^{\infty}a_{2n}\,k^{2n}$ are extracted by Lagrange inversion:
@@
Equation (10) provides, to our knowledge, novel closed-form expressions for the exact CE transport coefficients at all orders, determined entirely by the equilibrium velocity moments (2m1)!!(2m{-}1)!!. The explicit evaluation of the first few coefficients yields $a_{2}=-1,a_{4}=1,a_{6}=-4,a_{8}=27,a_{10}=-248,a_{12}=2830,a_{14}=-38232$, matching the sequence reported in Ref. [12]. As shown in Fig. 1, the coefficients grow factorially and the ratio $r_{n}=|a_{2(n+1)}/a_{2n}|$ increases roughly linearly; we later show that $r_{n}\sim 2n$, implying zero radius of convergence for the spatial CE series.
@@
The factorial divergence of the CE coefficients originates from the nature of the series $F(x)=\sum_{m=1}^{\infty}(-1)^{m}(2m{-}1)!!\,x^{m}$. The classical approach to analyzing $F$ is through its Borel transform, $\mathcal{B}[F]$. Using the identities $(2m{-}1)!!/m!=\binom{2m}{m}/2^{m}$ and $\sum_{m=0}^{\infty}\binom{2m}{m}y^{m}=(1-4y)^{-1/2}$, we obtain
@@
To understand the analytic structure of $\mathcal{B}[F]$, we note that $F$ is recovered from its Borel transform via the directional Laplace transform: $F(x)=\int_{0}^{+\infty}e^{-\sigma/x}\mathcal{B}[F](\sigma)\,d\sigma$. Since the Borel transform has a unique singularity on the negative real axis at $\sigma_{\star}=-1/2$, the Laplace contour along the positive real axis is entirely unobstructed. Consequently, the CE series is strictly Borel summable (exhibiting no Stokes phenomenon), and the Borel sum unambiguously reconstructs the exact dispersion relation $\hat{\omega}(k)$ without the need for lateral contour deformation or transseries completion.
Furthermore, the large-order growth of the CE coefficients follows the asymptotic behavior $a_{2n}\sim(-1)^{n}\,n!\,2^{n}$, rigorously explaining the ratio $r_{n}\sim 2n$ observed in Fig. 1.
@@
This strict summability stands in contrast to the temporal gradient expansion of Bjorken flow. In Ref. [10] the Borel transform of the Müller-Israel-Stewart temporal gradient series possesses its leading singularity at $\xi_{0}=3/(2C_{\tau\Pi})>0$ on the positive real axis, corresponding to the decay rate of the non-hydrodynamic mode. In that case, the Laplace contour is obstructed, rendering the series non-Borel summable and requiring a full transseries completion.
@@
In the spatial case, the divergence originates from an ultraviolet microscopic effect, namely the unbounded velocity tail of the equilibrium distribution. This is a fundamentally different mechanism from the temporal case, where the divergence is due to a macroscopic non-hydrodynamic mode. This difference explains both the negative sign, as the alternating CE coefficients reflect the competition between even and odd velocity moments, and the strict Borel summability. The Burnett instability is therefore not a fundamental pathology but a truncation artifact of this alternating divergent series, fully resolved by Borel resummation.
@@
Figure 2 confirms this: the Borel–Padé resummation of the CE series exactly reconstructs the non-perturbative attractor, while the spectral polynomial branches $P_{n}=0$ converge slowly to the hydrodynamic manifold with increasing truncation order. By contrast, the first-order CE truncation (classical diffusion), $\mathrm{CE^{2}}:\hat{\omega}=-k^{2}$, and the second-order Burnett-type approximation, $\mathrm{CE^{4}}:\hat{\omega}=-k^{2}+k^{4}$, diverge at moderate $k$.
@@
Since the factorial divergence originates from the unbounded velocity tail, it is tempting to consider a relativistic kinetic equation, imposing relativistic causality, resulting in the velocity phase space being bounded by the speed of light.
@@
Taken together with the results of Refs. [10] and [14], our findings suggest a unifying picture: the hydrodynamic gradient expansion, whether temporal or spatial, is always Borel summable, and its resummation invariably reconstructs the unique non-perturbative attractor defined by the (slow) invariant manifold.
Our results suggest that the rigorous passage from kinetics to hydrodynamics, a central aspect of Hilbert’s sixth problem [9], need not rely on a convergent perturbative expansion. Instead, we conjecture that hydrodynamics can be systematically derived from kinetic theory through Borel resummation of the factorially divergent gradient series, providing a non-perturbative route to the hydrodynamic limit.

### 03 Apr 2026 v2

@@ Abstract
Far-from-equilibrium kinetic systems collapse onto a hydrodynamic attractor, traditionally approximated by a gradient expansion. While temporal gradient series are non-Borel summable and require transseries completions, the analytic structure of the spatial expansion has remained elusive. Here, we derive exact closed-form Chapman–Enskog coefficients at all orders via Lagrange inversion and prove that the non-relativistic spatial gradient series, though factorially divergent, is strictly Borel summable. Furthermore, we show that this divergence originates from unbounded Galilean velocities; enforcing relativistic causality yields a convergent spatial hydrodynamic expansion with finite radius. Together with prior temporal results, our findings suggest that the hydrodynamic gradient expansion is always Borel summable, pointing to a non-perturbative route from kinetic theory to hydrodynamics.
@@
The classical treatment of non-equilibrium systems in kinetic theory relies on a systematic gradient expansion of the distribution function, known as the Chapman–Enskog (CE) expansion [6], truncated at finite order to yield macroscopic fluid dynamics. The validity of this expansion is controlled by the Knudsen number, the ratio of microscopic kinetic scales to the macroscopic scales of the coarse-grained hydrodynamic description. In relativistic systems, gradients of macroscopic fields normalized by the local temperature play an analogous role. However, it is well established that the resulting gradient expansion is purely asymptotic and diverges factorially [5].
@@
Nevertheless, numerical [7] and experimental [1] evidence has shown that far-from-equilibrium systems can still be quantitatively described by hydrodynamics, pointing to the existence of a hydrodynamic attractor.
@@
This phenomenon was first explored for temporal gradient expansions by Heller and Spaliński [11]. They demonstrated that for the highly symmetric, longitudinally expanding quark-gluon plasma (Bjorken flow) within Müller-Israel-Stewart framework, the fluid dynamics reduce to a set of ordinary differential equations. They showed that the attractor rigorously exists and its exact reconstruction requires generalized Borel resummation; the series is non-Borel summable along the positive real axis due to singularities corresponding to the decay of non-hydrodynamic modes, necessitating a full transseries completion.
@@
This attractor paradigm was subsequently generalized by Romatschke [18], who showed that conformal Bjorken flow across three distinct microscopic theories: (a) a variant of second order BRSSS hydrodynamics [3], (b) Boltzmann theory, and (c) strongly coupled $\mathcal{N}=4$ supersymmetric Yang Mills (via AdS/CFT), universally collapses onto a single attractor.
@@
These foundational studies focused exclusively on the temporal (longitudinal) gradient expansion, establishing that its divergence universally requires generalized Borel resummation. The spatial gradient expansion poses a distinct challenge: finite-order truncations develop short-wavelength instabilities, suggesting that a non-perturbative treatment may be required.
For the one-dimensional BGK kinetic equation, Karlin et al. [13] derived the exact invariance equation for the hydrodynamic manifold and constructed its solution through an iterative numerical procedure. A complementary non-perturbative approach was developed by Kogelbauer [15] in the framework of spectral theory: the hydrodynamic manifold was identified as the slow eigenmode of the associated Jacobi operator, yielding the exact diffusion eigenvalue and a critical wavenumber beyond which the slow mode ceases to exist. More recently, it was established [14] that the CE series is locally equivalent to this spectral closure to all orders in the Knudsen number, yet diverges factorially everywhere except at equilibrium.
What has remained open is whether the divergent spatial CE series nonetheless encodes the exact hydrodynamic manifold and can be resummed to reconstruct it.
@@
In this letter, we show that for non-relativistic Boltzmann kinetic theory, the spatial gradient expansion is factorially divergent but strictly Borel summable, uniquely reconstructing the spatial attractor. Also we show that once relativistic causality is imposed on the velocity space, the divergence is cured entirely, yielding a strictly convergent series with a finite radius.
@@
Starting from the one-dimensional kinetic Bhatnagar-Gross-Krook (BGK) [4] equation and setting the particle mass $m$, thermal energy $k_{B}T$, and relaxation time $\tau$ to unity, we write the non-dimensionalized kinetic equation for the distribution function $f(x,v,t)$ as
@@
In Fourier space this means $\hat{Z}_{\text{re}}=\hat{\Theta}_{\text{re}}(\lambda,k^{2})\,\hat{n}$ and $\hat{Z}_{\text{im}}=ik\,\hat{\Theta}_{\text{im}}(\lambda,k^{2})\,\hat{n}$, where the factor ikik in $\hat{Z}_{\text{im}}$ reflects the odd parity of the flux. The central requirement is the dynamic invariance condition: the macroscopic and microscopic time derivatives must agree exactly on this manifold [10, 12], bridging the kinetic and hydrodynamic descriptions in the spirit of Hilbert’s sixth problem.
@@
For small $k$ (equivalently large AA), the resolvent is expanded as a geometric series, $\frac{A}{A+v^{2}}=\sum_{m=0}^{\infty}(-1)^{m}(v^{2}/A)^{m}$, and integrated term by term using the Gaussian moments $\langle v^{2m}\rangle=(2m{-}1)!!$ :
@@
where $x=k^{2}(1{+}\hat{\omega})^{-2}=A^{-1}$. Since $x$ itself depends on $\hat{\omega}$, Eq. (Exact Chapman-Enskog Coefficients.—) is an implicit relation for $\hat{\omega}(k)$. The Chapman–Enskog coefficients $\hat{\omega}=\sum_{n=1}^{\infty}a_{2n}\,k^{2n}$ are extracted by Lagrange inversion:
@@
Equation (10) provides, to our knowledge, novel closed-form expressions for the exact CE transport coefficients at all orders, determined entirely by the equilibrium velocity moments (2m1)!!(2m{-}1)!!. The explicit evaluation of the first few coefficients yields $a_{2}=-1,a_{4}=1,a_{6}=-4,a_{8}=27,a_{10}=-248,a_{12}=2830,a_{14}=-38232$, matching the sequence reported in Ref. [13]. Using the asymptotic expansion of the exact slow diffusion eigenvalue, Ref. [14] found the same sequence of coefficients (up to an alternating sign), bridging their similarity to the sequence A000699 in Sloane’s Online Encyclopedia of Integer Sequences [16].
As shown in Fig. 2, the coefficients grow factorially and the ratio $r_{n}=|a_{2(n+1)}/a_{2n}|$ increases roughly linearly; we later show that $r_{n}\sim 2n$, implying zero radius of convergence for the spatial CE series.
@@
The factorial divergence of the CE coefficients can be traced to the (2m1)!!(2m-1)!! coefficient in the generating function $F(x)=\sum_{m=1}^{\infty}(-1)^{m}(2m{-}1)!!\,x^{m}$ appearing in (Exact Chapman-Enskog Coefficients.—), which implies zero radius of convergence for $F$. The analytic structure of $F$ can be revealed by the Borel transform. Using the identities $(2m{-}1)!!/m!=\binom{2m}{m}/2^{m}$ and $\sum_{m=0}^{\infty}\binom{2m}{m}y^{m}=(1-4y)^{-1/2}$, we obtain
@@
The original series $F$ can be recovered from its Borel transform via the directional Laplace integral,
$F(x)=\frac{1}{x}\int_{0}^{\infty}e^{-\sigma/x}\,\mathcal{B}[F](\sigma)\,d\sigma.$
@@
$\mathcal{B}[F](\sigma)$ has a unique singularity at $\sigma_{\star}=-1/2$ on the negative real axis, which means that the Laplace integration contour along $\sigma\in[0,\infty)$ is entirely unobstructed; $F$ is therefore strictly Borel summable, with no Stokes phenomenon and no ambiguity requiring lateral contour deformation.
Since the CE series $\hat{\omega}(k)=\sum a_{2n}\,k^{2n}$ is obtained from $F$ through the implicit relation (Exact Chapman-Enskog Coefficients.—) via Lagrange inversion (10), the Borel-plane singularity structure is inherited: the Borel transform of the CE series, $\mathcal{B}[\hat{\omega}](\sigma)=\sum_{n=1}^{\infty}\frac{a_{2n}}{n!}\,\sigma^{n}$, shares the same dominant singularity at $\sigma_{\star}=-1/2$.
Near this singularity, $\mathcal{B}[\hat{\omega}](\sigma)\sim C\,(\sigma-\sigma_{\star})^{-1/2}$; by Darboux’s theorem, the Taylor coefficients of a function with an algebraic singularity $(\sigma-\sigma_{\star})^{-\alpha}$ scale as $\sigma^{n}]\sim\sigma_{\star}^{-n}\,n^{\alpha-1}/\Gamma(\alpha)$. Setting $\alpha=1/2$ and $\sigma_{\star}=-1/2$ gives $a_{2n}/n!\sim(-2)^{n}\,n^{-1/2}/\sqrt{\pi}$, which yields
$a_{2n}\sim(-1)^{n}\,n!\,2^{n},$ (11)
explaining the linearly growing ratio $r_{n}=|a_{2(n+1)}/a_{2n}|\sim 2n$ observed in Fig. 2. The CE series is likewise strictly Borel summable: the physical dispersion relation $\hat{\omega}(k)$ is recovered without ambiguity from the Laplace integral
$\hat{\omega}(k)=\int_{0}^{\infty}\frac{1}{k^{2}}\,e^{-\sigma/k^{2}}\,\mathcal{B}[\hat{\omega}](\sigma)\,d\sigma,$
with no need for transseries completion.
@@
This strict summability stands in sharp contrast to the temporal gradient expansion of Bjorken flow. In Ref. [11] the Borel transform of the corresponding temporal gradient series possesses its leading singularity at $\xi_{0}=3/(2C_{\tau\Pi})>0$ on the positive real axis, corresponding to the decay rate of the non-hydrodynamic mode. In that case, the Laplace contour is obstructed, rendering the series non-Borel summable and requiring a full transseries completion.
@@
In the spatial case, the divergence originates from an ultraviolet microscopic effect, namely the unbounded velocity tail of the equilibrium distribution. This is a fundamentally different mechanism from the temporal case, where the divergence is due to a macroscopic non-hydrodynamic mode. This difference explains both the alternating sign pattern, arising from the geometric series structure of the velocity-space resolvent, and the strict Borel summability. The Burnett instability is therefore not a fundamental pathology but a truncation artifact of this alternating divergent series, fully resolved by Borel resummation.
@@
Figure 1 confirms this: the Borel–Padé resummation of the CE series exactly reconstructs the non-perturbative attractor, while the spectral polynomial branches $P_{n}=0$ converge to the hydrodynamic manifold with increasing truncation order. This is consistent with the spectral closure of Refs. [15, 14], where the exact diffusion eigenvalue was derived up to a critical wavenumber $k_{\text{crit}}=\sqrt{\pi/2}/\tau\approx 1.25$, beyond which the hydrodynamic mode merges with the essential spectrum. Our Borel resummation reconstructs the hydrodynamic attractor for $0<k<k_{\text{crit}}$ and analytically continues it beyond $k_{\text{crit}}$, where the hydrodynamic mode ceases to exist as an isolated spectral point. By contrast, the first-order CE truncation (classical diffusion), $\mathrm{CE^{(2)}}:\hat{\omega}=-k^{2}$, and the second-order Burnett-type approximation, $\mathrm{CE^{(4)}}:\hat{\omega}=-k^{2}+k^{4}$, diverge at moderate $k$.
@@
Since the factorial divergence originates from the unbounded velocity tail, it is tempting to consider a relativistic kinetic equation, where causality bounds the velocity phase space by the speed of light.
@@
Taken together with the results of Refs. [11] and [18], our findings point to a unifying picture: the hydrodynamic gradient expansion, whether temporal or spatial, is always Borel summable, and its resummation always reconstructs the unique non-perturbative attractor defined by the (slow) invariant manifold.
Our results indicate that the rigorous passage from kinetics to hydrodynamics, a central aspect of Hilbert’s sixth problem [10], need not rely on a convergent perturbative expansion. Instead, we conjecture that hydrodynamics can be systematically derived from kinetic theory through Borel resummation of the factorially divergent gradient series, providing a non-perturbative route to the hydrodynamic limit.
@@
Acknowledgments.—
Author thanks C. E. Frouzakis for comments on the initial draft.